OpenStudy (anonymous):

A polynomial function can be written as (x + 2)(x + 3)(x − 5). What are the x-intercepts of the graph of this function

OpenStudy (hybrik):

To find this, set this to (x + 2)(x + 3)(x − 5)=0 Find the numbers that can multiply to 0 to make 0=0 x= -2,-3,5 (-2+2 = 0) (-3+3 = 0) (5-5 = 0) So the x intercepts are -2,-3,5